### v2

# Beyond Fixed Rounds: Data-Free Early Stopping for Practical Federated Learning

Abstract
Federated Learning (FL) facilitates decentralized collaborative learning without transmitting raw data. However, reliance on fixed global rounds or validation data for hyperparameter tuning hinders practical deployment by incurring high computational costs and privacy risks. To address this, we propose a data-free early stopping framework that determines the optimal stopping point by monitoring the task vector’s growth rate using solely server-side parameters. The numerical results on skin lesion/blood cell classification demonstrate that our approach is comparable to validation-based early stopping across various state-of-the-art FL methods. In particular, the proposed framework requires an average of 45/12 (skin lesion/blood cell) additional rounds to achieve over 12.3%/8.9% higher performance than early stopping based on validation data. To the best of our knowledge, this is the first work to propose an data-free early stopping framework for FL methods.
footnotetext: Corresponding author
1 Introduction
Deep learning has driven significant advancements in medical imaging, utilizing large-scale datasets to achieve remarkable diagnostic performance [31]. However, the deployment of AI is strictly limited by stringent privacy regulations that prohibit the centralization of sensitive patient data [25, 27]. To overcome this barrier, Federated Learning (FL) has emerged as a promising decentralized paradigm that facilitates collaborative learning without transmitting raw data [23]. By ensuring that data remain localized at their source, FL preserves data sovereignty and strictly adheres to institutional governance and ethical standards [20]. Moreover, FL enables robust and generalizable learning in medical AI by effectively leveraging cross-institutional data diversity [3]. Thus, FL provides a scalable and privacy-preserving solution well suited for the secure development of collaborative medical AI systems [10].
Recent FL methods have evolved to enhance convergence stability and performance by refining optimization techniques at the client or server level [4]. The foundational approaches are typically grounded in Stochastic Gradient Descent (SGD), where FL methods like FedAvg, FedProx [21], SCAFFOLD [12], and FedDyn [1] regulate updates to stabilize the learning process. Subsequently, advanced FL methods have shifted towards Sharpness-Aware Minimization (SAM) [8], which seeks flat minima rather than merely minimizing loss values. This paradigm has led to the development of SAM-based FL methods—including FedSAM [26], FedSpeed [29], FedSMOO [28], FedGamma [6], FedLESAM [7], and FedWMSAM [22]—that smooth and stabilize the optimization trajectory.
Figure 1: Illustration of resource inefficiency in FL hyperparameter tuning. Since standard FL protocols use a fixed number of global rounds, ‘bad’ configurations waste computational and communication resources.
Despite the demonstrated effectiveness of recent FL methods, a practical constraint remains in their reliance on a pre-defined number of global rounds for training [14], as illustrated in Fig. 1. Similar to centralized learning, achieving optimal performance in FL requires extensive tuning of hyperparameters unique to each FL method or standard settings such as client-side learning rates [16]. In practice, this tuning process inevitably involves running ‘bad’ configurations that fail to converge to a satisfactory solution [13]. Under fixed-round protocols of FL, these ineffective configurations can cause high computational and communication overhead [18].
In this work, we propose a novel data-free early stopping framework that determines when to stop training using only the global model parameters at the server. Note that, unlike existing approaches relying on validation signals [15, 24], our framework adopts a purely model-driven stopping criterion. By avoiding the need for validation data, our framework strictly adheres to the FL paradigm of model-only transmission [11]. We show that our approach seamlessly integrates with 10 state-of-the-art FL methods and remains robust across medical imaging datasets. Moreover, our framework maintains consistent stability under various non-IID distributions, effectively handling data heterogeneity. The extensive experiments validate that our framework achieves generalization performance comparable to that of approaches relying on real validation data [33, 35].
2 Problem and Method
Figure 2: Illustration of the proposed data-free early stopping framework. The server monitors the growth rate of the task vector using only global model parameters and stops training once the growth rate falls below the threshold. Here, the color intensity of the task vector reflects its increasing magnitude, and vice versa.
2.1 Federated Setting
We consider a federated network comprising a central server and $N$ clients, designed for medical image classification tasks. The system aims to optimize the global model parameters $\theta$ by minimizing the global objective function $F(\theta)\triangleq\frac{1}{N}\sum_{n=1}^{N}F_{n}(\theta)$, where $F_{n}(\theta)$ denotes the local objective function calculated over the private dataset of the $n$-th client. In each global round, a subset of $M$ clients is randomly sampled to perform local learning. To simulate the non-IID nature, we consider three non-IID partitioning types following [19]: Label skew (Dirichlet, Pathological) and Quantity skew. For all considered distributions, the specific data allocation across clients is determined by a coefficient $c$.
2.2 Proposed Framework: Data-Free Early Stopping
We propose a novel early stopping framework that links federated convergence dynamics [5] with task vector characteristics [17] using only server-side global parameters, without relying on validation data. In detail, the FL process starts at global round $r=1$ with an initialized model $\boldsymbol{\theta}_{0}$. At each round $r\geq 1$, the central server aggregates $M$ client-side local updates to obtain the global model $\boldsymbol{\theta}_{r}$ through server-side optimization. We define the global task vector $\mathbf{v}_{r}\in\mathbb{R}^{d}$ as the cumulative displacement from the initialization:
$\mathbf{v}_{r}:=\boldsymbol{\theta}_{r}-\boldsymbol{\theta}_{0}=\sum_{k=1}^{r}\left(\boldsymbol{\theta}_{k}-\boldsymbol{\theta}_{k-1}\right).$ (1)
As training progresses, the global model moves away from the initialization, inducing an increasing task-specific displacement in the parameter space [9]. From an optimization perspective, each global update can be interpreted as a fine-tuning step resulting from the coupled dynamics of ClientOpt($\cdot$) and ServerOpt($\cdot$). Under standard smoothness assumptions, the task vector approximates the accumulated gradient flow as:
$\mathbf{v}_{r}\approx-\sum_{k=1}^{r}\gamma_{k}\nabla F(\boldsymbol{\theta}_{k-1}).$ (2)
Here, $\gamma_{k}$ represents the effective step size determined by the local learning rate, the number of local steps, and the aggregation scaling. Since FL satisfies stationarity conditions, i.e., $\lim_{r\to\infty}\|\nabla F(\boldsymbol{\theta}_{r})\|_{2}=0$, the growth of $\mathbf{v}_{r}$ diminishes as training converges. Thus, the accumulated optimization distance, $\delta_{r}:=\|\mathbf{v}_{r}\|_{2}$, gradually converges to a stable value.
To capture this convergence behavior, we introduce the growth rate $g_{r}$, which quantifies the relative increase of the accumulated distance:
$g_{r}=\frac{\delta_{r}-\delta_{r-1}}{\delta_{r-1}},\quad r\geq 2.$ (3)
As the learning trajectory stabilizes, $g_{r}$ tends to decrease, indicating that later global updates contribute marginally to the overall displacement. This behavior reflects the onset of convergence in the parameter space. To be consistent with data-based early stopping, the proposed criterion is restricted to two hyperparameters: a sensitivity threshold $\tau$ and a patience parameter $\rho$. In particular, we define a recursive saturation counter $\kappa_{r}$ as:
$\kappa_{r}=\mathbb{I}(g_{r}<\tau)\cdot(\kappa_{r-1}+1),\quad\kappa_{1}=0,$ (4)
where $\mathbb{I}(\cdot)$ denotes the indicator function. The federated training process is stopped at the round $r^{*}$ satisfying:
$r^{*}=\min\{r\geq 2\mid\kappa_{r}\geq\rho\}.$ (5)
The overall framework and procedure are shown in Fig. 2 and Algorithm 1.
Algorithm 1 Proposed Data-Free Early Stopping Framework
1:for $r=1$ to $R$ do
2:  for all selected clients in parallel do
3:   $\boldsymbol{\theta}_{r}^{m}\leftarrow\textsc{ClientOpt}(\boldsymbol{\theta}_{r-1})$
4:  end for
5:  $\boldsymbol{\theta}_{r}\leftarrow\textsc{ServerOpt}(\{\boldsymbol{\theta}_{r}^{m}\}^{M}_{m=1})$
6:  $r^{\star}\leftarrow\textsc{CheckEarlyStop}(\boldsymbol{\theta}_{r},r)$
7:  if $r^{\star}>0$ then return $\boldsymbol{\theta}_{r^{\star}}$
8:  end if
9:end for
10:return $\boldsymbol{\theta}_{R}$
[…]
13:Compute $\mathbf{v}_{r}$ using Eq. (1)
14:if $r\geq 2$ then
15:  Compute $g_{r}$ using Eq. (3)
16:  Update $\kappa_{r}$ using Eq. (4)
17:  if $\kappa_{r}\geq\rho$ then return $r$
18:  end if
19:end if
20:return 0
3 Experiment and Results
Figure 3: Test accuracy (%) and growth rate trajectories over global rounds with $\tau=0.01$ and $\rho=10$ for both proposed and validation-based early stopping. The zoom-in boxes highlight the regions around the early stopping points.
3.1 Experiment Setting
We evaluate the proposed approach against validation-based early stopping using validation loss or accuracy on skin lesion [30] and blood cell [2] image classification tasks. Note that the baselines use both training and validation splits in [34], while our approach uses only the training subset. All clients employ ConvNeXtV2 [32] as the local AI model. We benchmark with recent FL methods, including FedAvg, FedProx, SCAFFOLD, FedDyn, FedSAM, FedSpeed, FedSMOO, FedGamma, FedLESAM, and FedWMSAM. To simulate the federated setting, the dataset is partitioned across $N=100$ clients, and a subset of $M=10$ clients is randomly sampled for local training at each round. The experiments are repeated with $3$ random seeds, and our code is available at this repository.
3.2 Numerical Results
3.2.1 Effectiveness of Proposed Framework
To show the effectiveness of the proposed framework, as shown in Fig. 3, we evaluate FedAvg under label skew (Dirichlet) with $c=0.1$ and compare it against validation-based early stopping. On the skin lesion task, our approach stops, on average, +45+45 rounds later than validation-based early stopping, while achieving more than $12.3\%$ higher performance. For the blood cell task, training is extended by +12+12 rounds on average, yielding a mean performance gain of $8.9\%$. Moreover, the lower panels of Fig. 3 show that the proposed growth rate metric gradually decays over rounds, leading to stable stopping at the predefined threshold. Overall, the results indicate that the proposed early stopping framework achieves validation-level performance without requiring any validation data.
Table 1: Performance comparison under diverse non-IID data distributions and coefficient values $c$ with $\tau=0.01$ and $\rho=10$. Each value denotes the mean test accuracy difference between the proposed stopping point and the best-performing validation-based stopping point. The positive and negative values are highlighted in blue and red, respectively, where color intensity reflects the absolute magnitude of the difference. Note that bold and underlined values indicate the largest value for each $c$.
Method Label skew (Dirichlet) Label skew (Pathological) Quantity skew
$c=0.01$ $c=0.1$ $c=1.0$ $c=1$ $c=2$ $c=3$ $c=0.01$ $c=0.1$ $c=1.0$
Skin lesion task FedAvg +16.55 +3.42 +0.68 +3.02 +15.84 -2.98 +10.61 +4.08 -0.26
FedProx +12.62 +14.52 +2.93 +11.74 +10.50 +1.93 +10.47 +7.46 +5.29
FedDyn +20.87 +15.49 -2.40 +6.56 +13.06 +4.66 +18.07 +4.78 +8.78
SCAFFOLD +15.53 +1.92 +6.55 +7.53 +3.18 +6.11 -1.73 +3.07 +8.45
FedSAM +12.18 +13.44 -0.15 +2.50 +1.80 +3.56 +10.53 +5.67 +7.58
FedSpeed +24.82 +16.51 +0.60 +25.82 +16.94 +1.45 +24.94 +7.62 +9.43
FedSMOO +18.85 +17.60 +0.81 +24.06 +8.51 +2.34 +27.17 +5.06 +12.57
FedGamma +1.07 +4.64 +3.33 +2.76 +2.76 +3.78 +0.54 +0.78 +7.90
FedLESAM +19.65 +5.17 +1.36 -5.38 +3.38 +4.74 +15.15 +6.97 +3.98
FedWMSAM +15.73 +19.14 +17.81 +19.36 +29.59 +17.30 -0.23 -1.33 +3.78
Blood cell task FedAvg +7.69 -0.15 -1.34 -12.08 -1.26 +3.06 -0.23 -0.53 -0.40
FedProx +5.33 -0.20 -0.95 +4.37 +0.09 +0.94 +0.55 +0.06 -1.79
FedDyn +24.88 +5.49 +0.00 +27.70 +8.08 +1.90 +0.27 +0.37 +0.40
SCAFFOLD -4.14 -1.18 +1.31 +4.49 +0.83 +0.82 +1.08 +0.61 +3.42
FedSAM +7.91 +3.59 -1.21 +11.23 +3.71 +1.25 +0.85 +0.37 -0.32
FedSpeed +13.96 +4.90 +0.06 +37.16 +8.50 +1.10 -0.10 -0.12 +0.26
FedSMOO +16.08 +4.71 +0.56 +37.14 +8.60 +0.83 -0.05 -0.07 +0.39
FedGamma +2.93 -0.56 +0.51 +6.53 -1.52 +0.03 +6.08 -4.88 +3.18
FedLESAM -0.29 +2.77 -1.01 +7.42 +5.20 -1.13 -0.15 +0.17 -1.32
FedWMSAM -1.73 +15.77 -5.82 +16.89 -7.32 -5.33 -4.43 -0.22 -12.92
3.2.2 Impact of non-IID Data Distributions
To analyze the impact of non-IID data distributions, we evaluate the proposed framework under three representative data skew types across multiple $c$ values, as shown in Table 1. In detail, we analyze the performance differences at the respective stopping points between the proposed and validation-based early stopping approach. For the skin lesion task, the proposed approach achieves large average gains under severe heterogeneity ($c=0.01/1$), reaching approximately $15.8\%$, $9.8\%$, and $11.6\%$ for Dirichlet, Pathological, and quantity skew, respectively. As $c$ increases to $1.0/3$, the average gains decrease to around $3.2\%$, $4.3\%$, and $6.8\%$, indicating natural alignment with validation-based stopping as data distributions become IID.
For the blood cell task, a similar trend is observed, where the proposed framework yields average gains of approximately $7.3\%$, $14.1\%$, and $0.4\%$ under severe Dirichlet, Pathological, and quantity skew, respectively. Notably, across both tasks and skew types, the proposed approach achieves substantial gains under severe heterogeneity up to $29.6\%$$37.2\%$ (skin lesion/blood cell), which cannot be attributed to trivial decay of updates. This consistent pattern across tasks shows that the proposed criterion captures meaningful convergence under non-IID settings, rather than merely responding to diminishing updates. Thus, the proposed framework enables reliable hyperparameter tuning across diverse data distributions, with performance comparable to validation-based stopping.
Figure 4: Test accuracy (%) of FL methods evaluated under various threshold values with $\rho=10$ for both validation-based and proposed early stopping. The solid curves denote the mean test accuracy, while the shaded regions indicate the standard deviation.
3.2.3 Impact of Threshold ($\tau$)
To investigate the effect of the stopping threshold $\tau$ on the proposed framework, we vary $\tau\in\{0.005,0.01,0.05,0.1\}$ under label skew (Dirichlet) with $c=0.1$. Note that for validation-based early stopping, we report the best test accuracy obtained at the same $\tau$ using either validation loss or validation accuracy. As shown in Fig. 4, on the skin lesion task, the proposed framework shows a monotonic decrease in test accuracy as $\tau$ increases ($0.005\rightarrow 0.1$), with average reductions of approximately $19.4\%$ across FL methods. In particular, FedDyn decreases from $61.0\%$ and $63.9\%$ at $\tau\in\{0.005,0.01\}$ to $49.6\%$ and $45.8\%$ at $\tau\in\{0.05,0.1\}$, while FedSMOO decreases from $64.2\%$ to $46.3\%$ ($0.01\rightarrow 0.1$). This trend indicates that larger $\tau$ values cause early termination, which enables rapid screening of configurations but limits performance.
A similar trend is observed on the blood cell task when compared with validation-based early stopping. At $\tau\in\{0.005,0.01\}$, the proposed framework generally matches or exceeds the validation-based baseline across most FL methods, for example achieving $97.2\%$ versus $92.1\%$ for FedSpeed and $86.0\%$ versus $86.1\%$ for FedAvg. As $\tau$ increases to $0.05$ and $0.1$, the proposed accuracy drops markedly below the validation-based results, reaching $81.1\%$ and $74.6\%$ for FedDyn, and $80.9\%$ and $73.5\%$ for FedSMOO. This comparison shows that large $\tau$ values enable fast evaluation but stop training too early to reach the optimum round. By contrast, small $\tau$ values allow longer optimization, driving the model toward solutions closer to the optimal global model. Taken together, these results establish $\tau$ as a simple and effective control knob for balancing fast evaluation and convergence toward near-optimal models.
Table 2: Performance and stopping behavior of the proposed approach in terms of the stopping round ($r^{*}$) and corresponding test accuracy (Acc.), under $\tau=0.1$ and $\rho=10$. The $\Delta_{Acc.}$ and $\Delta_{r}$ denote the accuracy gain and round difference, respectively, relative to the best validation-based baseline at its stopping point.
Method Skin lesion task Blood cell task
Acc. (%) $r^{*}$ $\Delta_{Acc.}$ $\Delta_{r}$ Acc. (%) $r^{*}$ $\Delta_{Acc.}$ $\Delta_{r}$
FedAvg 14.29 15.0 - +4.7 12.50 16.0 - +6.0
FedProx 14.29 14.7 - +4.7 12.50 15.0 - +4.7
FedDyn 14.29 19.0 - +9.0 12.50 19.0 - +9.0
SCAFFOLD 14.29 26.0 - +16.0 12.50 22.3 - +12.3
FedSAM 14.29 15.0 - +5.0 12.50 14.0 - +4.0
FedSpeed 14.29 18.0 - +8.0 12.50 19.3 - +9.3
FedSMOO 14.29 17.7 - +7.7 12.50 17.7 - +7.7
FedGamma 14.29 25.3 - +15.3 12.50 24.0 - +14.0
FedLESAM 14.29 15.0 - +5.0 12.50 15.7 - +5.7
FedWMSAM 14.29 21.0 -0.18 +10.7 12.50 21.0 - +8.3
  • Note. The symbol ‘-’ denotes no difference.
3.2.4 Ablation Study
To validate the framework’s efficiency, we conduct an ablation study focused on handling bad configurations. In this scenario, the global model fails to learn, achieves only random-guess level accuracy $14.3\%$$12.5\%$ (skin lesion/blood cell). As discussed in the threshold analysis, we use a large $\tau$ to enable fast evaluation and early termination. As shown in Table 2, the proposed framework requires only 441616 additional rounds compared to the best validation-based baseline across all FL methods, far below the fixed budget of $500$ rounds. On average, the proposed framework stops $\Delta_{r}\approx+9$ rounds later for the skin lesion task and $\Delta_{r}\approx+8$ rounds later for the blood cell task, relative to the best validation-based baseline. Note that this is less than $2\%$ of the fixed-round budget, enabling rapid screening of bad configurations with minimal overhead. Therefore, the numerical results demonstrate that large $\tau$ enables efficient resource savings, particularly during early-stage tuning in FL.
4 Conclusion
In this work, we propose a data-free early stopping framework that identifies the stopping point via global task vector dynamics. The numerical results show that by tuning the threshold, our framework can either extend training for better performance or match the efficiency of validation-based baselines. Moreover, our framework significantly reduces the computational waste of fixed-round training by screening ineffective trials. The proposed approach stops comparably to validation-based approach, requiring only 9/8 (skin lesion/blood cell) additional rounds on average for FL methods. Thus, this work validates the feasibility of data-free early stopping, enabling resource-efficient tuning of FL methods.